# Evaluating LLMs’ Reasoning Over Ordered Procedural Steps

### v1

@@ 1 Introduction
Understanding and generating correctly ordered action sequences is a key aspect of reasoning. Many real world tasks, such as cooking recipes or carrying out technical procedures, require steps be completed in a precise order to achieve the intended outcome. LLMs have demonstrated strong performance on various reasoning tasks including arithmetic computation (Imani et al., 2023; Ahn et al., 2024), commonsense inference (Rajani et al., 2019), and question answering (Robinson et al., 2022). While much prior work has evaluated LLMs on step-by-step reasoning, their ability to reason over and reconstruct ordered procedural steps remains relatively underexplored.
@@
Step ordering tasks, where the correctness of the output depends on recovering a globally coherent sequence, pose a unique challenge. Most existing research focuses on predicting the immediate next step (Yong et al., 2025; Wang et al., 2023), rather than reconstructing the full sequence from a shuffled set. Moreover, prior evaluations rely only on accuracy (Quan and Liu, 2024), measuring exact matches between predicted and reference positions. This limits our ability to fully understand LLMs procedural reasoning. In this work, we evaluate LLMs’ step ordering capabilities using a curated dataset of food recipes due to their clearly defined structure and strong ordering constraints. As illustrated in Figure 1, the model receives a shuffled list of recipe instructions and must recover the correct sequence that reflects the intended preparation process. We incorporate complementary metrics—Kendall’s Tau for rank correlation, Normalized Longest Common Subsequence (NLCS) for subsequence preservation, and Normalized Edit Distance (NED) for reordering cost to provide a deeper analysis of model performance. We conduct a systematic evaluation across multiple LLMs under 0-shot and few-shot settings. We further analyze performance as a function of sequence complexity, examining how models respond to longer recipes and greater amounts of step shuffling. Our main contributions are:
  • We evaluate step-order reasoning in LLMs using a structured cooking recipe dataset for full sequence reconstruction.
  • We propose a multi-metric evaluation framework capturing partial correctness, subsequence alignment, and reordering cost.
  • We analyze performance variations with step count and shuffling difficulty, revealing gaps and challenges in procedural reasoning.
@@ 4 Dataset
Majumder et al. (2019) has introduced a dataset containing 230K recipes from Food.com [^1] https://www.food.com. From this corpus, we select 5,000 samples with 6 to 8 steps and 5 to 6 ingredients, ensuring moderate sequence length and complexity. Food recipes are inherently sequential, and prior work has treated step ordering as critical to successful execution (Wang et al., 2023). However, some recipes may have some steps that may be interchangeable without affecting the outcome (e.g., cutting onions and cutting potatoes). To focus on sequences where step order is necessary, we apply an additional curation step using a LLM to filter recipes requiring strict ordering (see Appendix A) which yields to 1,740 recipes. Each recipe provides a coherent step sequence $S=\{s_{1},\ldots,s_{n}\}$, which we shuffle randomly (with fixed seed) to produce $\hat{S}$. The task is to recover the original order from $\hat{S}$. We generate a permutation label $\pi\in\{1,\ldots,n\}^{n}$, where $\pi_{i}$ denotes the original position of the $i$-th step in the shuffled sequence. The dataset is balanced with $29.6\%$ (515 samples) having 6 steps, $36.7\%$ (638 samples) with 7 steps, and $33.7\%$ (587 samples) with 8 steps. We quantify the extent of step permutation by measuring the average positional displacement, defined as the mean absolute difference between the original position $p_{i}$ and the shuffled position $s_{i}$ of each step $i$ in a sequence of length $n$, i.e., $\frac{1}{n}\sum_{i=1}^{n}|p_{i}-s_{i}|$. This metric captures the average magnitude of step movement caused by shuffling. As shown in Figure 2, longer sequences show higher average displacement, increasing from 1.97 for 6-step to 2.62 for 8-step recipes, indicating greater complexity in step rearrangements. The median displacement remains at 2 across all lengths.
@@ 5.1 Inference Settings
We evaluate four instruction-tuned LLMs: Llama-3.1-8B-Instruct(Grattafiori et al., 2024), Mistral-7B-Instruct(Jiang et al., 2024), Gemma-2-9b-it(Team et al., 2024), and GPT-4o-mini(OpenAI et al., 2024) under three settings: zero-shot, 3-shot, and 5-shot. Model receives task instructions, the recipe name, and a list of shuffled steps (see Appendix B, C). It is expected to output the reordered steps and their corresponding indices. From 1,740 samples, we use 1,700 for testing and the rest for few-shot demonstrations.
@@ 5.2 Evaluation Metrics
We report accuracy at the step level: $\text{Acc}=\frac{1}{n}\sum_{i=1}^{n}(\hat{\pi}_{i}=\pi_{i})$. This metric measures the fraction of steps placed at the correct positions provides an measure of how often the model recovers the exact step location.
@@
Kendall’s tau is a rank correlation metric  (Lapata, 2006) that evaluates the relative order of all possible step pairs between the predicted permutation $\hat{\pi}$ and the ground truth $\pi$. It is computed as $\tau=\frac{C-D}{\frac{1}{2}n(n-1)}$, where $C$ is the number of concordant pairs and $D$ is the number of discordant pairs. It is suitable for assessing whether the predicted step sequence agrees with the ground truth in terms of relative step precedence, regardless of their absolute positions.
@@
Edit distance counts the number of insertions, deletions, or swaps required to convert the predicted order into the gold sequence. We use its normalized form (Marzal and Vidal, 2002), $\text{NED}=\frac{\text{EditDistance}(\hat{\pi},\pi)}{n}$. This metric measures the total transformation cost and is particularly sensitive to local misplacements.
@@
We compute the length of the longest common subsequence (LCS) between $\hat{\pi}$ and $\pi$, normalized by the length of the reference: $\text{NLCS}=\frac{\text{LCS}(\hat{\pi},\pi)}{n}$. This metric rewards the preservation of correct subsequences and reflects the extent to which a model recovers partial ordering structure. It is robust to small local reorderings and has been widely used in structured sequence evaluation.
@@
Table 1: Performance of different models across few-shot settings (0, 3, 5) using Accuracy (Acc), Normalized Longest Common Subsequence (NLCS), Kendall Tau (KTau), and Normalized Edit Distance (NED). Best values are bolded (lowest for NED).
@@ 6.1 Performance in Zero-Shot and Few-Shot Settings
Table 1 reports LLMs performance in 0-shot and few settings. All models show notable improvements from 0-shot to 3-shot prompting, with no gains beyond 3-shot. This suggests that a small number of demonstrations helps models learn structural reordering patterns, but further they do not add additional value. GPT-4o consistently achieves the best performance across all metrics. In the 3-shot setting, it reaches the highest accuracy (0.64), NLCS (0.85), and KTau (0.90), and the lowest NED (0.27), indicating better absolute positioning, strong preservation of subsequences, and minimal local reordering. Gemma-2 performs competitively whereas Mistral and Llama-3.1 fall behind across all metrics, often producing more fragmented sequences (lower NLCS) and higher reordering costs (higher NED), despite moderate KTau scores.
KTau values show that even when models make positional errors, they may still preserve correct relative ordering. For example, Llama-3.1 in 3-shot achieves 0.83 KTau despite only 0.45 accuracy, indicating good understanding of step precedence even with absolute misplacements. NED values further expose models’ tendency to make local misorderings— GPT-4o and Gemma-2B consistently yield the lowest values. NLCS emphasizes preservation of long subsequences; GPT-4o and Gemma-2B again score highest, indicating better retention of step continuity. Despite overall improvements in few-shot settings, all models exhibit gaps in fine-grained step-level reasoning.
@@ 6.2 Impact of Number of Steps on Ordering Performance
We analyze model performance by the number of steps in the sequence (n), where longer sequences indicate increased complexity. As shown in Figure 3, with n increasing from 6 to 8, a general performance decline is observed across all models, reflecting the added difficulty in recovering longer step sequences. GPT-4o consistently performs best, maintaining high accuracy (0.73 $\rightarrow$ 0.56), strong subsequence alignment (NLCS: 0.88 $\rightarrow$ 0.82), and low edit cost (NED: 0.21 $\rightarrow$ 0.31) as complexity increases. Gemma-2B shows similar robustness, with slightly lower performance. Llama-3.1 and Mistral performance drops more significantly indicating their tendency to produce fragmented and disordered outputs under increased complexity.
@@ 6.3 Impact of Step Average Positional Displacement on Model Performance
We further assess robustness to reordering by analyzing model performance with respect to average positional displacement. As shown in Figure 4 with displacement increasing, indicating more severe shuffling, all models show performance drops. For Accuracy, Gemma starts highest but drops sharply from near 1.0 to  0.4, while GPT-4 declines more gradually, indicating greater stability. Mistral performs lowest overall but declines steadily, suggesting consistent underperformance rather than heightened sensitivity. In NLCS, all models degrade with displacement, but GPT-4 and Gemma maintain relatively stable and higher scores. NED increases with displacement, reflecting greater divergence from the reference; here, GPT-4 shows a smaller increase compared to the other models.
@@ 7 Conclusion
We evaluated four LLMs on step ordering tasks using four complementary metrics. All models improved from 0-shot to 3-shot prompting, with no gains beyond. GPT-4o consistently achieved the best performance, followed by Gemma-2, while Llama and Mistral performed less reliably. As sequence length and reordering complexity increased, performance declined across the board. While models often preserved relative ordering (high KTau) and subsequences (high NLCS), they still struggled with precise step level reasoning highlighting limitations in LLMs’ procedural understanding.
@@ A.2 Dataset Curation Examples
Some examples of data curation is provided in Figure 6.
Figure 6: Examples of data curation
@@ B.1 Prompt
The prompt used for evaluation of the LLMs for step ordering is given in Figure 7.
@@ Appendix C Experimental Details
The table shows the hyperparameters of the LLM model used for experimentation and their respec- tive values. We have used 1 A100 GPUs.

### v2

@@ 1 Introduction
Understanding and generating correctly ordered action sequences is a key aspect of reasoning. Many real world tasks, such as cooking recipes or carrying out technical procedures, require steps be completed in a precise order to achieve the intended outcome. LLMs have demonstrated strong performance on various reasoning tasks including arithmetic computation (Imani et al., 2023; Ahn et al., 2024), commonsense inference (Rajani et al., 2019), question answering (Robinson et al., 2022; Anika et al., 2025), and multimodal reasoning and understanding tasks (Miah et al., 2025, 2023). While much prior work has evaluated LLMs on step-by-step reasoning, their ability to reason over and reconstruct ordered procedural steps remains relatively underexplored.
@@
Step ordering tasks, where the correctness of the output depends on recovering a globally coherent sequence, pose a unique challenge. Most existing research focuses on predicting the immediate next step (Yong et al., 2025; Wang et al., 2023), rather than reconstructing the full sequence from a shuffled set. Moreover, prior evaluations rely only on accuracy (Quan and Liu, 2024), measuring exact matches between predicted and reference positions. This limits our ability to fully understand LLMs procedural reasoning. In this work, we evaluate LLMs’ step ordering capabilities using a curated dataset of food recipes due to their clearly defined structure and strong ordering constraints. As illustrated in Figure 1, the model receives a shuffled list of recipe instructions and must recover the correct sequence that reflects the intended preparation process. We use complementary metrics, including Kendall’s Tau to measure rank correlation, Normalized Longest Common Subsequence (NLCS) to assess subsequence preservation, and Normalized Edit Distance (NED) to quantify reordering cost, providing a deeper analysis of model performance. We conduct a systematic evaluation across multiple LLMs under 0-shot and few-shot settings. We further analyze performance as a function of sequence complexity, examining how models respond to longer recipes and greater amounts of step shuffling. Our main contributions are:
  • We present a comprehensive evaluation of step-order reasoning abilities in LLMs using a structured cooking recipe dataset, going beyond next-step prediction to assess full sequence reconstruction.
  • We introduce a multi-metric evaluation framework that captures partial correctness, subsequence alignment, and reordering cost—offering a richer picture of model behavior than accuracy alone.
  • We analyze how model performance varies with step count and shuffling difficulty, revealing performance gaps and highlighting ongoing challenges in LLM’s procedural reasoning
@@ 4 Dataset
Majumder et al. (2019) has introduced a dataset containing 230K recipes from Food.com [^1] https://www.food.com. From this corpus, we select 5,000 samples with 6 to 8 steps and 5 to 6 ingredients, ensuring moderate sequence length and complexity. Food recipes are inherently sequential, and prior work has treated step ordering as critical to successful execution (Wang et al., 2023). However, some recipes may have some steps that may be interchangeable without affecting the outcome (e.g., cutting onions and cutting potatoes). To focus on sequences where step order is necessary, we apply an additional curation step using a LLM to filter recipes requiring strict ordering (see Appendix A) which yields to 1,740 recipes. Each recipe provides a coherent step sequence $S=\{s_{1},\ldots,s_{n}\}$, which we shuffle randomly (with fixed seed) to produce $\hat{S}$. The task is to recover the original order from $\hat{S}$. We generate a permutation label $\pi\in\{1,\ldots,n\}^{n}$, where $\pi_{i}$ denotes the original position of the $i$-th step in the shuffled sequence.
The dataset is balanced with $29.6\%$ (515 samples) having 6 steps, $36.7\%$ (638 samples) with 7 steps, and $33.7\%$ (587 samples) with 8 steps.
We quantify the extent of step permutation by measuring the average positional displacement, defined as the mean absolute difference between the original position $p_{i}$ and the shuffled position $s_{i}$ of each step $i$ in a sequence of length $n$, i.e.,
$\frac{1}{n}\sum_{i=1}^{n}|p_{i}-s_{i}|$
This metric captures the average magnitude of step movement caused by shuffling. Figure 2 visualizes the displacement distribution, highlighting mean values and variability. Longer sequences show higher average displacement, increasing from 1.97 for 6-step to 2.62 for 8-step recipes, indicating greater complexity in step rearrangements. The median displacement remains at 2 across all lengths.
@@ 5.1 Inference Settings
We evaluate five instruction-tuned LLMs: Llama-3.1-8B-Instruct(Grattafiori et al., 2024), Mistral-7B-Instruct(Jiang et al., 2024), Gemma-2-9b-it(Team et al., 2024), GPT-4o-mini(OpenAI et al., 2024) and Qwen3-8b(Team, 2025) under three settings: zero-shot, 3-shot, and 5-shot.
Each model is prompted with a set of task instructions, the recipe name, and a list of shuffled procedural steps (see Appendix B, C ). The model is expected to output both the reordered step sequence and the corresponding order as a list of indices. Qwen3-8b is a reasoning model, and thinking mode was enabled during evaluation. From the 1,740 samples, we use 1,700 as the test set and remaining samples for few-shot demonstrations.
@@ 5.2 Evaluation Metrics
We report accuracy at the step level:
$\text{Acc}=\frac{1}{n}\sum_{i=1}^{n}(\hat{\pi}_{i}=\pi_{i})$
This metric measures the fraction of steps placed at the correct positions and provides a measure of how often the model recovers the exact step location.
@@
Kendall’s tau is a rank correlation metric  (Lapata, 2006) that evaluates the relative order of all possible step pairs between the predicted permutation $\hat{\pi}$ and the ground truth $\pi$. It is computed as
$\tau=\frac{C-D}{\frac{1}{2}n(n-1)}$
where $C$ is the number of concordant pairs and $D$ is the number of discordant pairs. It is suitable for assessing whether the predicted step sequence agrees with the ground truth in terms of relative step precedence, regardless of their absolute positions. This metric is sensitive to pairwise inversions and captures global ordering consistency.
@@
Edit distance counts the number of insertions, deletions, or swaps required to convert the predicted order into the gold sequence. We use its normalized form (Marzal and Vidal, 2002),
$\text{NED}=\frac{\text{EditDistance}(\hat{\pi},\pi)}{n}$
This metric measures the total transformation cost and is particularly sensitive to local misplacements. NED is an error-based metric; lower values indicate better sequence similarity.
@@
We compute the length of the longest common subsequence (LCS) between $\hat{\pi}$ and $\pi$, normalized by the length of the reference:
$\text{NLCS}=\frac{\text{LCS}(\hat{\pi},\pi)}{n}$
This metric rewards the preservation of correct subsequences and reflects the extent to which a model recovers partial ordering structure. It is robust to small local reorderings and has been widely used in structured sequence evaluation.
Together, these metrics capture global, local, and partial structural alignment between predicted and target step sequences.
@@
Qwen3 0-shot 0.71 0.88 0.92 0.22
3-shot 0.63 0.82 0.88 0.30
5-shot 0.62 0.81 0.87 0.30
Table 1: Performance of different models across few-shot settings (0, 3, 5) using Accuracy (Acc), Normalized Longest Common Subsequence (NLCS), Kendall Tau (KTau), and Normalized Edit Distance (NED). The best and second-best results across all models are highlighted (lowest for NED).
@@ 6.1 Performance in Zero-Shot and Few-Shot Settings
Table 1 reports LLMs’ performance in 0-shot and few-shot settings. Most models, i.e., Llama-3.1, Mistral, Gemma-2, and GPT-4o, show notable improvements from 0-shot to 3-shot prompting, whereas Qwen3 maintains strong performance even without demonstrations. This suggests that a small number of demonstrations helps models learn structural reordering patterns. However, across all models, performance plateaus beyond 3-shot as no model shows meaningful gains with five examples, indicating limited additional value from further demonstrations.
Qwen3 achieves the best overall performance across all metrics in the 0-shot setting, reaching the highest accuracy (0.71), NLCS (0.88), and KTau (0.92), and the lowest NED (0.22), indicating strong intrinsic capabilities for accurate and consistent reordering than other models. This superior performance suggests that Qwen3’s reasoning abilities allow it to better understand and model the logical structure of sequences, enabling it to maintain correct absolute positions and preserve subsequences effectively. GPT-4o ranks second overall, performing best among the remaining models in the 3-shot setting, with high accuracy (0.64), NLCS (0.85), and KTau (0.90), and low NED (0.27), indicating better absolute positioning, strong preservation of subsequences, and minimal local reordering. Gemma-2 performs competitively whereas Mistral and Llama-3.1 fall behind across all metrics, often producing more fragmented sequences (lower NLCS) and higher reordering costs (higher NED), despite moderate KTau scores.
KTau values show that even when models make positional errors, they may still preserve correct relative ordering. For example, Llama-3.1 in 3-shot achieves 0.83 KTau despite only 0.45 accuracy, indicating good understanding of step precedence even with absolute misplacements. NED values further expose models’ tendency to make local misorderings, with Qwen3 achieving the lowest scores, followed by GPT-4o and Gemma-2, indicating minimal local reordering. NLCS emphasizes preservation of long subsequences; again, Qwen3 attains the highest score, demonstrating stronger retention of step continuity compared to other models. Despite these improvements, all models still exhibit gaps in fine-grained step-level reasoning, suggesting remaining challenges in capturing detailed procedural structure.
@@ 6.2 Impact of Number of Steps on Ordering Performance
We analyze model performance in the three-shot setting by the number of steps in the sequence (n), where longer sequences indicate increased complexity. As shown in Figure 3, with n increasing from 6 to 8, a general performance decline is observed across all models, reflecting the added difficulty in recovering longer step sequences. GPT-4 consistently performs best, maintaining high accuracy (0.73 $\rightarrow$ 0.56), strong subsequence alignment (NLCS: 0.88 $\rightarrow$ 0.82), and low edit cost (NED: 0.21 $\rightarrow$ 0.31) as complexity increases. Gemma 2B shows similar robustness, with slightly lower performance. Qwen 3 performs competitively, surpassing LLaMA 3.1 and Mistral on most metrics, with stable subsequence alignment (NLCS $\approx$ 0.86 $\rightarrow$ 0.79) and consistent ranking correlation (KTau $\approx$ 0.89 $\rightarrow$ 0.86). In contrast, LLaMA 3.1 and Mistral show more pronounced declines across all metrics, indicating a stronger tendency to produce fragmented and disordered outputs under increased complexity.
@@ 6.3 Impact of Step Average Positional Displacement on Model Performance
We further assess robustness to reordering by analyzing model performance with respect to average positional displacement. As shown in Figure 4, as displacement increases, indicating more severe shuffling, all models exhibit noticeable performance degradation across metrics. For Accuracy, Gemma-2 starts highest but drops sharply from near 1.0 to approximately 0.4, while GPT-4 declines more gradually, demonstrating greater stability. Qwen3 achieves strong initial performance and remains competitive with GPT-4, showing moderate declines in accuracy and NLCS as displacement increases. In contrast, LLaMA 3.1 remains consistently lower, and Mistral performs worst overall, maintaining a steady but low trajectory across all displacement levels. For NLCS, GPT-4, Gemma-2, and Qwen3 maintain relatively high subsequence alignment compared to LLaMA 3.1 and Mistral. KTau follows a similar trend, with GPT-4 and Qwen3 sustaining high ranking correlation and smaller declines under increased displacement. NED rises with displacement, reflecting larger deviations from the reference ordering, where GPT-4 and Qwen show smaller increases compared to the other models, indicating stronger resistance to disorder.
@@ 7 Conclusion
We evaluated five LLMs on step ordering tasks using four complementary metrics. Most models improved from 0-shot to 3-shot prompting, with no additional gains from five examples, whereas Qwen3 maintained strong performance in the 0-shot setting but degraded in the 3-shot setting. No gains were observed from 3-shot to 5-shot for any model. Qwen3 achieved the best overall performance, followed by GPT-4o and Gemma-2, while Llama-3.1 and Mistral were less reliable. Performance declined as sequence length and reordering complexity increased. Although models often preserved relative ordering (high KTau) and subsequences (high NLCS), they continued to struggle with precise step-level reasoning, highlighting limitations in LLMs’ procedural understanding.
@@ A.2 Dataset Curation Examples
Some examples of data curation are provided in Figure 6. The examples demonstrate which food recipes was selected by the LLMs along with its reasoning.
@@ B.1 Prompt
The prompt used for evaluation of the LLMs for step ordering is given in Figure 7. This is the prompt for 0-shot setting. Demonstrations were incorporated for 3-shot and 5-shot settings.
Figure 5: LLM prompt used for data curation
B.2 Few-Shot Examples Selections
Figure 6: Examples of data curation
@@ Appendix C Experimental Details
The table shows the hyperparameters of the LLM models used for experimentation and their respective values. We used 1 A100 GPU for all experiments. For Qwen3-8B, the maximum number of generated tokens was set to 2048, and thinking mode was enabled, whereas for all other models ‘max_new_tokens‘ was 512.